An economy is described by: AS = 10p, C = 80 + .8(Y – .1Y – 5), I = 4, G0= 35 and NX = 5 – p – 0.12Y. Parliament enacts a rule requiring a balanced budget b2 = 0 and mandates a reduction in G2 < 35. With this reduction in spending, National Output will be equal to
Answer #1

Demand side y: C+I+G - NX y 8o+0.8 (Y-0.14-5) + 4 + 35+5-2 Y = $ 124 + 0.724 - 4 - P Y-0.724 120P Y 120-P 0.28Given, As. lop At Equilibrium, AS Y lop 120-R 0-28 120-P = 120 218p- 3. 8p jp= 31.57 Before change y: 31 5.1 [4. lopTC = 80 +0.8(4-0.14-5) = 80 +0.724-4 BY= -C OT c (consump I-C to efficient) 1) LAG l-c Go - 35 G2 < 35 Lets say G2-34 roa =D4= Nowe, Due to 1 1-0e72 X 1 8G) Loy AY: 3.57D4= Nowe, Due to 1 1-0e72 X 1 8G) Loy AY: 3.57Balanced budget is achieved by keeping the same amount of change in G and T.

As Balanced Budget= G - T

Therefore, ∆G = ∆T

I have assumed that ∆G = 1 = ∆T

Now, Balanced Budget multiplier is given by adding the individual multipliers of G and T.

Which is nothing but 1

Because, [1/(1 - c)] + [ -c / (1 - c)] = 1

So, change is national income will be always be equal to change in G or T.

∆Y = ∆G = ∆T

New national product = Y + ∆G

Here, with an example of ∆G=1, new national product is 316.7.